# Vision foundation models: can they be applied to astrophysics data?

### v3

@@ 5 Conclusions
Our results demonstrate that state-of-the-art vision foundation models can immediately perform classification of optical galaxies and source detection of radio galaxies with 72 - 88% precision. In the case of optical galaxy classification, all foundation models out-performed supervised training. F1 scores of close to 0.85 for 800 labeled training images, 10% of the available total, show that foundation models can be used to great effect for initial investigations into using machine learning for scientific tasks. Of the models tested, Vision Transformers DINOv2 and MSN performed the best, along with CNN ResNet50. MAE, despite its success on natural images, is optimized for reconstruction loss and failed to encode characteristics which are highly relevant for galaxy classification.
Shared characteristics of the model training distribution and the images of GMNIST might have contributed to good classification performance. Even though GMNIST images contain a lot of empty space, a good portion of the image is occupied by the galaxy of interest. If stars and other objects are present in the background, they are small relative to the galaxy, even if the galaxy is seen edge-on. On the other hand, classification of radio galaxies was challenging, as even the main radio source was small and the images were dominated by noise and patterns from reconstruction. In this case, supervised training outperformed all foundation models except MSN.
For a dataset like RGZ with a clear distribution shift from training to down-stream task data, applicability of foundation models is limited. The models did retain some relevant information about the data – in this case, mainly the information about bright peaks – and designing the data augmentations or label schemes to take advantage of that could improve results.
The need for compatibility between the dataset, down-stream task, and foundation model appears in the source detection example as well. Once images were scaled so that the Vision Transformer patch size was smaller than the average radio source, source detection results were encouraging. High mAP@50 scores showed that foundation models larger than ResNet 18 could identify radio sources, especially when the backbone was allowed to fine-tune. However, improvements due to fine-tuning were inconsistent and (in the case of MSN) much smaller than might be expected. This situation might improve given more optimal methods of fine-tuning.
Even in cases where results are good, such as source detection, there remains a large gap in performance between tasks done on scientific images and tasks done on natural images. State-of-the-art classification on thousand-class ImageNet-1k has a top-1 accuracy of 92%, while the best object detection on the standard COCO dataset has a mAP@50 of 0.73, which may seem low in comparison to our results but is the average metric across all 90 classes. Closing the gap in performance is necessary to achieve scientifically meaningful results, but it will have to be done with far less computing and human resources than are available to large companies. Techniques which can operate with very few labeled training examples and by training a small number of network parameters will be especially valuable for applying vision foundation models to scientific images. In this respect, the semi-supervised learning proposed in [47] is a very interesting option for further investigation.
In this study, we considered very simple way of fine-tuning. Perhaps a more powerful fine-tuning method, based on the simultaneous use of labeled and unlabeled data, might enhance the results, as suggested in [47]. This would be particularly exciting for astrophysics, where unlabeled data is plentiful. In future work, we will examine the potential of such semi-supervised learning framework for fine-tuning.

### v4

@@ 6 Conclusions
Our results demonstrate that state-of-the-art vision foundation models can immediately perform classification of optical galaxies and source detection of radio galaxies with F1 scores of 0.72 - 0.88. In the case of optical galaxy classification, all foundation models out-performed supervised training. F1 scores close to 0.87 with only 800 labeled training images (10% of the total dataset) show that foundation models can be used to great effect for early-stage machine learning applications in scientific contexts. Of the models tested, SigLIP and AM-RADIO, pre-trained on multi-modal datasets, delivered the best performance, with Vision Transformers such as DINOv2 and MSN also excelling over ResNet-50, which is used or compared with in much of the astrophysics literature. Only MAE, which is optimized for reconstruction loss, failed to encode many task-relevant features for galaxy classification.
Dataset characteristics strongly influenced model performance. For GMNIST, the alignment between the pre-training distribution and the task dataset likely contributed to the strong classification results, as galaxies occupied a significant portion of the image with minimal interference from noise or unrelated objects. Conversely, RGZ images, with small central sources and images dominated by noise and patterns from reconstruction, presented a clear distribution shift that limited the applicability of foundation models. In this case, supervised training outperformed all foundation models except SigLIP and AM-RADIO. While models retained some relevant features, such as bright peaks, we speculated that information about the flux islands, also relevant to classification, was lost due to the original pre-processing. Adapting dataset-specific pre-processing or label schemes could improve results, though significant domain expertise remains essential for working with radio astronomy datasets.
Source detection results further underscored the importance of aligning data to model architectures and tasks. When radio galaxy images were resized to ensure that transformer patch sizes were slightly smaller than average source dimensions, performance improved markedly. We saw that ViTs did well to detect compact sources in MGCLS, while ResNet was better at finding RGZ galaxies. Models pre-trained on multi-modal data usually did not perform as well as those pre-trained on only images, especially when freezing the backbone. Additionally, fine-tuning improvements were inconsistent and (in the case of MSN on RGZ) much smaller than might be expected. This situation might improve given a more optimal detection head architecture, or additional hyperparameter tuning.
Despite promising results, there remains a significant gap between the performance of foundation models on scientific datasets and their state-of-the-art results on natural image benchmarks. For example, ImageNet-1k classification achieves a top-1 accuracy of 92%, and COCO object detection achieves an average mAP@50 of 0.73 across 90 classes. Closing this performance gap will require techniques that maximize results from limited labeled data and computational resources, a challenge particularly relevant in domains like astrophysics. Semi-supervised learning frameworks, such as those proposed by Voloshynovskiy et al. (2020), offer exciting possibilities to make use of the abundance of unlabeled data to enhance fine-tuning methods.
In this study, we employed simple fine-tuning strategies, which, while effective, leave room for improvement. Future work will investigate advanced fine-tuning methods that integrate labeled and unlabeled data, potentially enabling more robust and resource-efficient application of vision foundation models to astrophysical image analysis. This iterative and adaptive approach will be essential for bridging the gap between cutting-edge machine learning methods and the unique demands of astronomical research.
@@
Acknowledgements
This work has been done in partnership of the Swiss SKA consortium which is funded by the State Secretariat for Education, Research and Innovation (SERI).
MGCLS data products were provided by the South African Radio Astronomy Observatory and the MGCLS team and were derived from observations with the MeerKAT radio telescope. The MeerKAT telescope is operated by the South African Radio Astronomy Observatory, which is a facility of the National Research Foundation, an agency of the Department of Science and Innovation.
We would like to thank Philip Denzel for the RGZ object detection dataset.
OB and DP were supported by the SNF Sinergia grant CRSII5-193826 “AstroSignals: A New Window on the Universe, with the New Generation of Large Radio-Astronomy Facilities”.
MD and VK were supported by the SNF Sinergia project (CRSII5-193716), “Robust deep density models for high-energy particle physics and solar flare analysis (RODEM)”.
Appendix A Attention Visualization
Figure 6: Images for the following attention maps
Figure 7: Attention maps for all models, RGZ compact source (left, labeled in the dataset as 1C-1P) and extended source (right, labeled in the dataset as 2C-2P), shown for each model and each attention block.
Figure 8: Attention maps for all models, GMNIST smooth-and-round galaxy (left) and unbarred spiral (right), shown for each model and each attention block.
Figure 9: Attention maps for all models, MGCLS multi-source crop (left). ImageNet (right), shown for each model and each attention block.
We visualize the attention maps for sample images from each dataset, which are shown in Figure 6. One map per attention block is shown; ViT-Base, the backbone architecture for all Vision Transformers used in this study, has twelve blocks.
Maps of blocks with higher numbers, which typically capture high-level semantic relationships as opposed to earlier blocks which focus on more low-level, localized features, do not always display meaningful features when calculated for astrophysical images (Figures 7, 8 and 9 left side). This is especially evident with SigLIP and AM-RADIO. Models pre-trained on natural images only (DINOv2, MSN, and MAE), however, do display the typical behavior described above in attention maps calculated for an ImageNet example (Figure 9, right side). Examination of the attention maps therefore illustrates that the semantic relationships, both local and global, learned from natural image pre-training do not always have meaningful counterparts in astrophysical images.